Write the equation of the quadratic function with roots 0 and 2 and a vertex at (1, 5).


Answer by Alan3354(69443)

y = x*(x-2) gives the 2 roots, but the vertex is below the x-axis.
y = -x*(x-2) inverts it.
y = -x^2 + 2x gives the 2 roots, but the vertex is at 1,1.
Multiply it by 5:
y = -5x^2 + 10x works